### v1

# A Systematic Survey and Critical Review on Evaluating Large Language Models: Challenges, Limitations, and Recommendations

Abstract
Large Language Models (LLMs) have recently gained significant attention due to their remarkable capabilities in performing diverse tasks across various domains. However, a thorough evaluation of these models is crucial before deploying them in real-world applications to ensure they produce reliable performance. Despite the well-established importance of evaluating LLMs in the community, the complexity of the evaluation process has led to varied evaluation setups, causing inconsistencies in findings and interpretations. To address this, we systematically review the primary challenges and limitations causing these inconsistencies and unreliable evaluations in various steps of LLM evaluation. Based on our critical review, we present our perspectives and recommendations to ensure LLM evaluations are reproducible, reliable, and robust.
1 Introduction
The evolution of LLMs has transitioned from simple generative models predicting the next word to advanced systems capable of following instructions and solving complex problems Zhao et al. (2023a). Early models like GPT Radford et al. (2018) could generate coherent text but were limited to simple tasks, whereas instruction-tuned LLMs Ouyang et al. (2022); Chung et al. (2022) like ChatGPT111https://openai.com/index/chatgpt/ greatly enhanced their versatility and ability to execute specific commands. This shift has revolutionized the development of real-world applications powered by LLMs.
With the advancements and broad applicability of LLMs, it is essential to properly evaluate them to ensure they are safe to use. This is indeed important not only for academic benchmarks but also for business use cases. Consequently, understanding the bottlenecks of current evaluation methods, and developing strategies to address these challenges are crucial for standardizing evaluations and enabling reliable use of LLMs in practical applications. Nonetheless, evaluating LLMs is as complex and resource-intensive as their development, involving multiple levels or aspects.
Existing reviews Chang et al. (2024); Guo et al. (2023b); Zhuang et al. (2023); Minaee et al. (2024); Liang et al. (2022) related to the evaluation of LLMs often focus only on benchmark tasks, datasets, and evaluation criteria, neglecting the broader complexities. This oversight can undermine the reliability of evaluation by ignoring issues like robustness and reproducibility. While some recent studies Balloccu et al. (2024); Mao et al. (2023) have investigated data contamination Ravaut et al. (2024) and evaluation malpractices in LLM evaluation, their focus is limited to only assessing ChatGPT, overlooking other LLMs, as well as the entire evaluation pipeline.
More recently, Biderman et al. (2024) discussed the reproducibility problem in existing evaluations of LLMs and introduced a library to address this. However, their work lacks comprehensive discussions on how aspects like reliability or robustness impact LLM evaluation and how to address them. Hence, existing LLM evaluation studies often focus on individual aspects in a scattered manner, resulting in findings that are only sparsely useful.
To mitigate this gap, this paper brings together the discussions to address the fundamental challenges and limitations in LLM evaluations that emerge from diverse evaluation setups. First, we craft a schematic workflow of the evaluation pipeline in practical settings (presented in Section 2) for a systematic study. We then examine each step in the evaluation workflow, uncovering various inconsistencies and decision-making complexities affecting reproducibility, reliability, and robustness (see Section 3). Based on our findings, we provide a principled guideline in Section 4 to address current limitations in LLM evaluation.
2 Overview of LLM Evaluation Process
The following components are crucial for LLM evaluation: Evaluation Setup, Response Generation, and Evaluation Methodology Chang et al. (2024). Each component has its own challenges, which we discuss in Section 3. These components in an evaluation workflow are shown in Figure 1.
2.1 Evaluation Setup
Benchmark Selection:
To initiate the evaluation process of LLMs, the first step is selecting appropriate benchmarks. We categorize the benchmarking datasets into the following: general capability benchmarks, specialized benchmarks, and other diverse benchmarks. We refer to general capability benchmarks as the ones that are often used for evaluation upon the release of an LLM (e.g., MMLU Hendrycks et al. (2020b), HumanEval Chen et al. (2021)). In addition, there are specialized benchmarks that measure specific capabilities of LLMs (e.g., MT-Bench for chatting capabilities Zheng et al. (2024)). There are also other benchmarks that usually combine multiple benchmarks to evaluate LLMs on diverse task (e.g., HELM Liang et al. (2022)). We provide more details on each category in Appendix A.1.
Model Selection:
Selecting the appropriate model from the numerous LLMs currently available is crucial for ensuring a fair evaluation, as it helps to avoid risks such as data contamination and unfair comparisons. For a detailed discussion on prominent LLMs, see Appendix A.2.
2.2 Response Generation
Once the benchmarks and the models are selected, the next step in the evaluation process is to design the prompt and set up the decoding parameters for response generation. In the prompt design step, decisions on what type of prompting (e.g., zero-shot or few-shot) would be used are taken. Moreover, configuring the decoding parameters (e.g., temperature) is important to ensure optimal performance Shi et al. (2024). More discussions on this are provided in Appendix A.3 and A.4.
2.3 Evaluation Methodology
Parsing Script Design:
Evaluating LLM-generated responses is difficult because they often produce verbose outputs (see Table 4 for some examples). Therefore, parsing scripts are often necessary Laskar et al. (2023a); Jahan et al. (2024) to extract target labels before applying evaluation metrics, ensuring alignment with evaluation criteria to maintain reliability.
Evaluation Approach:
The evaluation approach can be divided into the following: automatic evaluation, human evaluation, LLMs as evaluators. In automatic evaluation, before applying task-specific metrics (e.g., F1, Exact Match, Perplexity Jelinek et al. (1977)), parsing scripts are often utilized to extract the targeted answer, especially in discriminative tasks. Human evaluation is required to ensure qualitative assessments of LLM responses (e.g., measuring clarity, coherence, factuality) van der Lee et al. (2021). Recently, human evaluation based on the Elo-based rating system Zheng et al. (2024) has gained a lot of attention. Since human evaluation is time-consuming, the utilization of LLMs as evaluators to assess other LLMs has become a popular evaluation approach Huang et al. (2024a); Chiang and Lee (2023). More details on LLM evaluation approaches are in Appendix A.6.1.
3 Challenges in Evaluating LLMs
We examine challenges and limitations in the evaluation process of LLMs based on three dimensions: reproducibility, reliability, and robustness.
3.1 Reproducibility
Reproducibility, the ability to consistently replicate model results under the same conditions, is a major challenge in generative models Biderman et al. (2024). The primary challenge is the lack of comprehensive documentation for each part of the evaluation cycle, including benchmarking datasets, prompt construction, model details, decoding strategy, response parsing, and evaluation methodology Kosch and Feger (2024); McIntosh et al. (2024). Table 1 presents an analysis by Balloccu et al. (2024), revealing that a relatively low percentage of the analyzed papers shared their resources. Below, we discuss factors impacting reproducibility in the evaluation step.
3.1.1 Missing Details on Data & Models Used
Availability (%) | Comparison (%) | ||||
Prompt | Code | Prompt + Code | Model Version | Fair | Unfair |
90.6 | 53.3 | 50.0 | 29.3 | 20.7 | 79.3 |
Benchmarking Data: One factor that can negatively impede the ability to reproduce results is not releasing the exact data used for evaluation Balloccu et al. (2024). Many studies evaluate LLMs on only a subset of existing datasets Bang et al. (2023); Kocoń et al. (2023), while others use the exact benchmarking datasets Laskar et al. (2023a); Qin et al. (2023). Despite the expectation not to compare results across studies using different subsets of the data, such comparisons often occur, as discussed by Balloccu et al. (2024). Nonetheless, without explaining the sampling strategy, or releasing the subsets used for evaluation (and possibly their responses), reproducing results using different data subsets of the same size is challenging.
Model Versions: The information regarding the version of a model being used is also missing in many studies Biderman et al. (2024); Balloccu et al. (2024), creating reproducibility concern (see Table 1). The continuous updates of the closed-source models, often with undisclosed changes can also impact reproducibility. With these updates, earlier versions are often deprecated, and results from these versions may not apply to newer models Chen et al. (2023b), making prior evaluation results to be no longer reproducible Laskar et al. (2023a); Bang et al. (2023); Qin et al. (2023); Kocoń et al. (2023). Therefore, it is crucial to specify the model versions used Biderman et al. (2024); Balloccu et al. (2024), while model owners should keep earlier versions available.
3.1.2 Lacking Response Generation Details
Prompting: The lack of details behind how the prompts are designed may make the findings in different literature inconsistent. For instace, variations in prompt design can lead to significantly different results, as seen in various studies Jahan et al. (2024); Laskar et al. (2023a); Qin et al. (2023); Bang et al. (2023). While few-shot learning is found to outperform zero-shot in the original evaluation conducted by the authors of various LLMs OpenAI (2023); Anil et al. (2023); Touvron et al. (2023b), many independent evaluations demonstrate that adding few-shot examples does not necessarily outperform zero-shot models in every task Ye et al. (2023a); Jahan et al. (2024). This raises the concern of whether certain prompt engineering techniques or optimizations to select few-shot samples were applied in the original evaluations. Hence, not disclosing the details behind how the prompt is designed or how the few-shot examples are selected can hinder reproducibility.
Decoding Strategy: LLMs are sensitive to decoding parameters, leading to significant performance variations based on the chosen settings Roziere et al. (2023); Touvron et al. (2023b). However, crucial details on their selection are excluded in existing literature OpenAI (2023); Team et al. (2023); Qin et al. (2023); Bang et al. (2023); Laskar et al. (2023a); Kocoń et al. (2023). This lack of transparency raises reproducibility concerns, which could be responsible for inconsistent results across studies even when similar prompts are used. For instance, Qin et al. (2023) found that adding output length restrictions in the prompt to generate summaries in no more than N words led to a performance drop in the SAMSum dataset Gliwa et al. (2019). However, Laskar et al. (2023a) found that such controlled experiments led to a gain in performance in the SAMSum dataset.
3.1.3 Evaluation Methods Unavailable
Parsing Scripts: LLM-generated responses often require parsing scripts to extract desired information. However, as demonstrated in Table 1, Balloccu et al. (2024) observed in their analysis that almost half of the LLM evaluation papers do not release any codes. We also observe that most studies (these include both the LLM technical reports, as well independent evaluations) do not release their parsing scripts OpenAI (2023); Team et al. (2024, 2023); Qin et al. (2023); Bang et al. (2023); Kocoń et al. (2023). Nonetheless, inaccurate design of parsing scripts may lead to different evaluation results Laskar et al. (2023a). Thus, the unavailability of parsing scripts would complicate result comparisons while impacting reproducibility Balloccu et al. (2024); Biderman et al. (2024).
Evaluation Approach: LLMs are increasingly used to evaluate other LLMs in development Zheng et al. (2024). Concerns arise due to the use of closed-source LLMs as evaluators, as their frequent updates can affect reproducibility Verga et al. (2024); Chen et al. (2023b). Moreover, Chen et al. (2023b) observed significant behavioral changes in closed-source LLMs over short periods. Such reproducibility concerns are also observed in prior research that used LLMs as evaluators. For instance, Chiang and Lee (2023); Zheng et al. (2024) found that using closed-source LLMs as the judge could collide with human evaluations, whereas Fu et al. (2023b) observed the opposite. Since the recently proposed Prometheus-2 Kim et al. (2024a) model is an open-source alternative and demonstrates a strong correlation with humans, utilizing open-source LLMs as the judge can help mitigate the reproducibility issues prevalent with closed-source LLMs.
3.2 Reliability
Reliability, the ability to trust that outcomes are as intended, is another challenge encountered during evaluation. Issues like contamination/inaccurate labels in the data, irrelevant evaluation methods, and unfair comparisons may impact the reliability of the findings, which we discuss below.
3.2.1 Data and Model Integrity Issues
Data Integrity: Errors in benchmarks undermine accurate conclusions and model comparisons, rendering evaluations of LLMs unreliable. An integrity-compromising factor is the presence of incorrect gold labels. For instance, existing issues in the gold labels of the widely used MMLU Hendrycks et al. (2020b) dataset have led to the development of MMLU-Pro Wang et al. (2024b) and MMLU-Redux Gema et al. (2024). Recently it was also found that the coding benchmarks, HumanEval Chen et al. (2021), lacked essential test cases, leading to the development of an advanced version, HumanEvalPlus Liu et al. (2024b).
Despite these improvements, many recent studies continue to use the older versions of datasets. For instance, despite the release of HumanEvalPlus, HumanEval is still used to benchmark LLM coding performance Team et al. (2023); Jiang et al. (2023); Li et al. (2023c); Roziere et al. (2023); Gloeckle et al. (2024); Team et al. (2024); Wong et al. (2023), potentially providing misleading insights. In addition, outdated labels in existing benchmarks undermine reliability of gold references. For example, in tasks like open-domain question answering, which demand real-world knowledge, many gold labels become outdated over time, as noted by Laskar et al. (2023a). Consequently, even if LLMs produce correct answers, comparing them to obsolete gold labels can yield inaccurate results. Moreover, in tasks like summarization, LLM-generated summaries are often favored over human-annotated gold references Pu et al. (2023); Zhang et al. (2024b); Ding et al. (2022).
Contamination in Existing Models: Contamination occurs when a benchmarking dataset is used in training, reducing result reliability and validity Sainz et al. (2023a); Zhou et al. (2023b); Shi et al. (2023). Ensuring benchmarking examples are excluded from training data is essential to maintain reliable results. Since LLMs are pre-trained on vast amounts of text data available on the internet, this could lead to unfair evaluations if LLMs have already encountered these datasets during their pre-training phase Ravaut et al. (2024); Xu et al. (2024); Balloccu et al. (2024).
Nonetheless, most prior LLM evaluation work focusing on zero-shot evaluation did not conduct any data contamination tests Laskar et al. (2023a); Bang et al. (2023); Qin et al. (2023); OpenAI (2023); Team et al. (2023), raising concerns about whether these evaluations truly represent the zero-shot capabilities of LLMs. Recent research has also demonstrated a strong possibility of data contamination in many datasets used to evaluate different LLMs Sainz et al. (2023b); Li and Flanigan (2023); Golchin and Surdeanu (2023); Ravaut et al. (2024); Xu et al. (2024); Zhang et al. (2024a); Oren et al. (2023); Balloccu et al. (2024). With the current generation of LLMs being extremely capable of learning new skills with minimal amounts of data, exposing them to evaluation data may undermine the measurement of their true capabilities. Since the possibility of data contamination has led to the development of new versions of existing datasets (e.g., utilizing GSM-8K to construct GSM-1K Zhang et al. (2024a)), it is crucial to use reliable and fair evaluation datasets.
3.2.2 Lack of Fairness by Manipulating Response Generation
Prompt Hacking:
One major concern in terms of lack of fairness in LLM evaluation is the possibility of prompt hacking Schulhoff et al. (2023), which involves manipulating input prompts to a language model to elicit desired responses (e.g., biasing the outputs, or taking unfair advantages by using specific few-shot examples). While the performance of LLMs depends on many factors relevant to how the prompt is structured, most work Bang et al. (2023); Qin et al. (2023); Laskar et al. (2023a), even the official technical reports Team et al. (2023); OpenAI (2023); Anthropic (2024) of different LLMs lack the necessary details behind prompt construction (e.g., missing scientific validity on why a certain prompt was preferred over others, how the few-shot examples are selected, etc.). This makes the claims regarding the effectiveness and limitations of certain LLMs in comparison to others questionable222https://crfm.stanford.edu/2024/05/01/helm-mmlu.html. Recognizing these parallels underscores the need for transparency and robust methodologies to ensure fairness in AI research and development.
Lack of Transparency in Decoding Parameters: Shi et al. (2024) demonstrated that extensive tuning of decoding parameters could improve the performance during inference. However, how the different decoding parameters are selected is often underexplored in existing evaluations Laskar et al. (2023b); OpenAI (2023); Team et al. (2023); Qin et al. (2023); Bang et al. (2023); Laskar et al. (2023a), as discussed in Section 3.1. This poses the risk of optimizing the parameters on test sets to improve performance.
3.2.3 Inappropriate Evaluation Methodology
Inaccurate Design of Parsing Scripts: As Laskar et al. (2023a) observed, evaluating LLMs entirely with an automated approach based on the answer extracted using parsing scripts may lead to an error of up to more than 10% difference in many tasks. This raises questions about the reliability of LLM evaluations that solely depend on parsing scripts without validating the scripts’ effectiveness for the task. To tackle this, Laskar et al. (2023a) proposed a hybrid approach combining parsing script-based automatic evaluation with human-in-the-loop Wu et al. (2022); Laskar et al. (2022a). Initially, the parsing script extracts answers from LLM-generated responses. If any issues arise, humans resolve them, enhancing the reliability of parsing-based automatic evaluation.
In Figure 2, we demonstrate the differences between automatic and hybrid evaluation in Open-Domain QA333NQ-Open Kwiatkowski et al. (2019), WebQuestions Talmor and Berant (2018), TriviaQA Joshi et al. (2017)) and reading comprehnesion datasets444SQuAD-V2 Rajpurkar et al. (2018), Race-High and Race-Middle Lai et al. (2017). The figure highlights the influence of human intervention on results in open-domain QA, where LLMs may generate synonymous or time-sensitive correct answers, potentially rendering gold answers outdated Laskar et al. (2023a). Parsing script-based automatic evaluation is found to be reliable in Race datasets for reading comprehension, whereas notable discrepancies are observed in the SQuAD-V2 dataset. Therefore, there’s a need for designing dependable parsing scripts and involving humans when appropriate.
Evaluation Approaches Lacking Relevancy: In generative tasks, utilizing automatic string-based matching techniques may not be reliable as well. For instance, Laskar et al. (2023a) observed that despite LLMs scoring quite poorly on the ROUGE metric compared to SOTA summarization models, humans often prefer LLM-generated responses. Moreover, recent research observed potential biases while using LLMs as evaluators, such as LLMs preferring responses generated by LLMs of the same series, positional bias Stureborg et al. (2024); Wu and Aji (2023); Wang et al. (2023b); Bai et al. (2024). To mitigate this, Verga et al. (2024) proposed a new technique that leveraged multiple LLMs as juries instead of using a single LLM as the judge. This approach demonstrates higher correlations with humans, while mitigating biases.
3.3 Robustness
While there are many evaluation benchmarks currently available, existing work mostly relies on evaluating LLMs on some common benchmarks, this raises the question of whether the performance of LLMs in these common benchmarks in existing settings reflects their true capabilities and limitations. In this section, we study the robustness of existing LLM evaluations.
3.3.1 Lacking Generalized Evaluation
Limiting Evaluation to Certain Scenarios: Interestingly, it has been observed in recent research that certain performance gains in a specific dataset may not necessarily imply that it would also improve the performance in other datasets for similar tasks Jahan et al. (2024); SambaNova (2024). For instance, Jahan et al. (2024) observes that not a single LLM has superiority over other LLMs across all biomedical datasets and tasks. This is also evident if we compare the results between LLaMA-3 and Qwen2 reported in Qwen2 (2024). As shown in Figure 3, while the Qwen2 model outperforms LLaMA-3 on most datasets, it falls short on GPQA and MBPP. Interestingly, for coding tasks, Qwen2 significantly outperforms LLaMA-3 on the HumanEval dataset Chen et al. (2021) but not on the MBPP dataset Austin et al. (2021). Meanwhile, existing common benchmarks also do not take into account some specific settings, such as how LLMs perform in long context scenarios, as recent research demonstrated that LLMs often struggle to generate the correct answer when relevant information does not appear at the beginning or end of the input context Liu et al. (2024c). This highlights the importance of evaluating the generalized performance of LLMs across a set of diverse benchmarks and settings,instead of limiting evaluation to only common benchmarks like MMLU Hendrycks et al. (2020b).
Tokenizer | Vocab | MMLU | MMLU-Pro | MixEval | MixEval-Hard |
LLaMA-2 | 32,000 | 0.52 | 0.45 | 0.29 | 0.11 |
LLaMA-3 | 128,256 | 0.27 | 0.21 | 0.09 | 0.03 |
Mistral | 32,000 | 0.59 | 0.51 | 0.31 | 0.11 |
Qwen2 | 151,646 | 0.22 | 0.17 | 0.08 | 0.02 |
Diversity and Coverage in Benchmarks: Although benchmarking datasets are designed to address specific problems and objectives, the variation and complexity of language within these datasets are often unclear. Liang et al. (2022) highlighted that better coverage in benchmarking datasets would enhance the comprehensiveness of the model’s evaluation. While different language models use different tokenizers to represent the benchmarking dataset, it also leads to variations in what is evaluated across models.
As can be seen in Table 2, we conducted a small-scale analysis for LLaMA-2 Touvron et al. (2023b), LLaMA-3,555https://llama.meta.com/llama3/ Mistral Jiang et al. (2023), and Qwen2666https://github.com/QwenLM/Qwen2 on two benchmarking datasets with varying complexities: MMLU Hendrycks et al. (2020b) and its more challenging version, MMLU-Pro Wang et al. (2024b), as well as MixEval Ni et al. (2024) and its harder version, MixEval-Hard. Our findings indicate that these datasets cover a relatively small portion of the model’s capabilities. Specifically, for MixEval, as the datasets became more diverse and dynamic, the vocabulary coverage for the tokenizer decreased. This trend continued as the datasets increased in difficulty, with vocabulary coverage further declining.
3.3.2 No Tuning of Prompt and Decoding Parameters
While various combinations of decoding parameters may lead to differences in results Shi et al. (2024), possibly due to high computing requirements, existing LLM evaluation work mostly undermines the necessity of evaluating how the model performance may vary depending on its variations. Similar to the absence of decoder parameter tuning, most prior work also evaluated LLMs using only a single prompt Laskar et al. (2023a); Qin et al. (2023); Bang et al. (2023); Kocoń et al. (2023); Jahan et al. (2024). However, in the real world, users express themselves with diverse word choices, varying semantics and syntaxes, alongside minor discrepancies (e.g., misspellings or differing punctuation styles). To further examine the effects of prompt variations, we conduct an experiment using GPT-4o (2024-04-09) and GPT-3.5-Turbo (0125) OpenAI (2023), as well as Claude-3-Opus (2024-02-29) Anthropic (2024) with the prompts used by Laskar et al. (2023a) and Qin et al. (2023) in the SAMSum dataset. For this experiment, the default parameters for respective LLMs are used.
As shown in Figure 4, the restricted prompting method by Laskar et al. (2023a) consistently outperforms the unrestricted approach across all three models. Conversely, the restricted prompting method by Qin et al. (2023) fails to surpass the unrestricted approach for GPT-3.5 and GPT-4o. However, it surprisingly outperforms the unrestricted method, indicating the significant impact of prompt tuning across models. Evaluating language models with a single prompt lacks fairness Zhu et al. (2023b), yet it remains common practice Bang et al. (2023); Qin et al. (2023); Laskar et al. (2023a). Minor prompt variations can lead to diverse outcomes for different models Alzahrani et al. (2024); An et al. (2023); Zhang et al. (2024a); Sclar et al. (2023); Lanham et al. (2023); Biderman et al. (2024), highlighting the need to compare benchmarks across multiple prompts. Using automated prompt tuning techniques like Meta Probing Agents Zhu et al. (2024) can ensure robustness to prompt variations.
3.3.3 Evaluation Method’s Generalizability and Correlation Shortcomings
While automatic evaluations are usually utilized in discriminative tasks, they may not be applicable to every task, as demonstrated by Jahan et al. (2024) that parsing scripts are not usable in certain discriminative tasks like relation extraction. Jahan et al. (2024) also noted a significant performance gap between the string-matching-based ROUGE metric Lin (2004) and the contextual similarity-based metric BERTScore Zhang et al. (2019) in text summarization. While larger models achieve better accuracy, they involve a speed-accuracy trade-off Parvez et al. (2019), leading to higher costs and latency Laskar et al. (2023b); Fu et al. (2024b). While metrics like perplexity are widely used to evaluate language models Chen et al. (2023c), Huang et al. (2024b) found that quantized LLaMA-3 versions have lower output confidence than the original. They noted similar model rankings for perplexity and a common-sense QA dataset. However, Hu et al. (2024) found no correlation between perplexity and long context understanding tasks, highlighting the need for robust evaluations with human-correlated metrics.
This raises another question, whether automated evaluations and LLM-as-a-judge correlate with human evaluations (e.g., Elo ratings). Zheng et al. (2024) demonstrated significant correlations between Elo ratings, LLM-as-a-judge, and automated evaluations. However, recent research Alzahrani et al. (2024) suggest that automated evaluations, especially those using multiple-choice questions, can yield unstable rankings with minor changes in evaluation methods. Given this instability, it prompts us to question why these automated tests should align with human Elo ratings despite demonstrating such inconsistencies. In our view, we should focus not only on correlating scores but also on how well a benchmark’s rankings align with the gold standards. Analysis in Table 3 for GPT-4 OpenAI (2023), Gemini Team et al. (2023), and Claude-3 Anthropic (2024) reveals two key observations: (i) MMLU rankings disagree with LMSys Chatbot Arena and (ii) MMLU rankings vary among themselves due to implementation differences.
Chatbot | HELM | Vellum | |
Model | Arena | MMLU | MMLU |
GPT-4o-2024-05-13 | 1 (1) | 2 (2) | 1 (1) |
GPT-4-Turbo-2024-04-09 | 5 (3) | 3 (3) | 3 (3) |
GPT-4-0125-preview | 6 (4) | 5 (5) | 4 (4) |
Gemini-1.5-Pro | 4 (2) | 4 (4) | 13 (6) |
Gemini-1.5-Flash | 10 (6) | 10 (6) | 10 (5) |
Claude-3-Opus-2024-02-29 | 7 (5) | 1 (1) | 2 (2) |
4 Recommendations and Best Practices
We’ve outlined the primary challenges in evaluating LLMs thus far. In light of these challenges, a crucial question arises: How can we enhance the evaluation process for LLMs? Crafting a structured framework that’s both practical and easy to implement is daunting, given the complexities of generative LLM development. Previous studies have tended to focus on specific evaluation aspects without offering comprehensive guidelines for the entire evaluation cycle, leaving researchers without clear guidance. Before diving into recommendations for each evaluation stage, it’s important to acknowledge three key factors shaping current LLM evaluation practices: inherent randomness in generative models, significant computational demands, and insufficient documentation across stages.
Evaluation Setup: Selecting benchmarks for model assessment is crucial. Rather than simply replicating past choices, researchers should align datasets with required capabilities. To ensure robustness, datasets should vary across expected LLM capabilities (e.g., long-context understanding), tasks (e.g., summarization), and language complexity (e.g., vocabulary coverage). Ideally, a metric should measure dataset diversity. For model selection, conduct contamination tests between the chosen model and benchmarks using relevant techniques Ravaut et al. (2024). This acts as an additional filter for benchmarking datasets, ensuring selection of unseen ones measuring intended capabilities. Meanwhile, for reproducibility, document any subset use of benchmarking datasets, along with the selected model version. In addition, throughout scientific history, intelligence progress has evolved across generations. Tests from a decade ago may appear simplistic compared to today’s standards (e.g., Math Olympiads, ICPC programming contests). Refreshing LLM evaluations periodically can effectively communicate standard capabilities in both open and closed-source LLM markets and ecosystems (e.g., chatbots, translation tools). Hence, to ensure reliability, verify if the dataset has updated versions and incorporate them if available (e.g., HumanEvalPlus Liu et al. (2024b), MMLU-Pro Wang et al. (2024b))
Response Generation: For reproducibility, thorough documentation of prompts and parameter settings is essential (e.g., specifying how few-shot samples are selected). To ensure reliability, it’s crucial to justify why specific prompts and parameter settings are chosen over others, and provide comparisons with alternative options. As for robustness, running experiments with diverse prompts and parameters is key to showcasing their effectiveness and limitations across different scenarios. In resource-constrained environments, conducting experiments with diverse evaluation settings may pose challenges, yet it remains vital to perform robust evaluations on at least a subset of samples.
Evaluation Methodology: To ensure reproducibility, the parsing scripts and the output data used for evaluation should be published. Meanwhile, sanity-checking on the parsing script should be done to ensure reliability and robustness of the designed parsing script. This can be done by creating test cases for various response types, and then verifying (with human intervention if possible) whether the parsing script can reliably extract the targeted answer from the generated response. Meanwhile, reliance on string-based metrics like ROUGE should be minimized in favor of qualitative evaluations to ensure the reliability of the chosen evaluation methodology. Given the cost and time constraints of human qualitative evaluation, LLM-based evaluators can be used as alternatives but must be validated for potential biases (e.g., multiple LLMs as juries instead of using a single LLM as the judge Zheng et al. (2024)). Finally, robust evaluation using task-specific metrics is encouraged with the metrics that lack alignments with humans should be avoided.
5 Conclusions and Future Work
In this paper, we systematically survey the challenges and limitations in evaluating LLMs. We identified significant inconsistencies and complexities at various stages of the evaluation pipeline, impacting the reproducibility, reliability, and robustness of the results. These issues underline the necessity for a standardized and systematic approach to evaluating LLMs to ensure their reliable usage in real-world applications. By comprehensively reviewing the current evaluation practices, we have provided a set of recommendations aimed at enhancing the consistency and reliability of LLM evaluations. Therefore, future work should focus on developing and adopting standardized evaluation protocols for LLMs to address the identified inconsistencies and complexities. This includes creating benchmark datasets, evaluation metrics, and proper documentation of the evaluation settings to ensure reproducibility, reliability, and robustness.
Limitations
One limitation of this work is that it is focused only on the evaluation phase of the LLM development cycle. Therefore, the challenges and limitations that happen during the training phase of LLMs are left out of the scope of this paper. Nonetheless, with the rapid growth of LLM technologies and huge financial incentives, it is essential to conduct a fair and reliable evaluation of LLM, alongside ensuring robustness and reproducibility, which is the focus of this work.
Another limitation of this study is that it does not study how to prevent closed-source LLMs from getting access to the online benchmarks. For instance, assume we have two entities: model developers and evaluators. Evaluators do not want to expose their data to the modeling team. Conversely, model developers do not want to release their model weights due to significant financial incentives. If evaluators use an API to get the responses, there is a risk that the queries may get exposed to the model developers. Therefore, without getting access to the weights, evaluators cannot reliably assess the models on their queries. Mathematically and technically, there is no fundamental way to solve this problem without altering the training dynamics which may not be an option for training teams.
Finally, the multimodal capability, in other words, the ability to understand both language and vision is another interesting capability of recently proposed LLMs Zhu et al. (2023a); Liu et al. (2024a, 2023b); Dai et al. (2024); Zhang et al. (2023); Ye et al. (2023b); Luo et al. (2024); Bai et al. (2023); Chen et al. (2023a). This has led to the development of many multi-modal benchmarks Fu et al. (2023a); Yu et al. (2023); Liu et al. (2023e); Li et al. (2023a, b); Liu et al. (2024a); Fu et al. (2024a); Lu et al. (2022); Guan et al. (2023); Li et al. (2023d); Qiu et al. (2024); Chen et al. (2024b). However, this paper was mostly focused on text-based NLP tasks and the evaluation of LLMs on multimodal benchmarks is left out for future work.
6 Ethics Statement
This paper only reviews the existing challenges and limitations in LLM evaluations and provides an opinion piece and recommendation to ensure reliable, robust, and reproducible evaluations of LLMs. Thus, this review does not pose any ethical concerns.
Appendix A Appendix
A.1 Benchmarking Datasets
General Capability Benchmarks:
To benchmark the performance of LLMs, researchers typically use a set of widely recognized datasets. These common benchmarks are employed by authors upon the release of an LLM to evaluate its general capabilities. One of the most frequently used benchmarks is the MMLU benchmark Hendrycks et al. (2020b), which assesses LLMs’ overall knowledge and reasoning abilities across various subjects. Other common benchmarks focus primarily on evaluating the commonsense reasoning capabilities of LLMs Wei et al. (2022a), such as HellaSwag Zellers et al. (2019), PIQA Bisk et al. (2020), SIQA, Sap et al. (2019), WinoGrande Sakaguchi et al. (2021), OpenBookQA Mihaylov et al. (2018), ARC Clark et al. (2018). In addition, the TruthfulQA dataset Lin et al. (2021) is used to measure the truthfulness of an LLM, while the TyDi QA dataset Clark et al. (2020) is used for evaluating the information seeking question answering capability across diverse languages. For assessing coding capabilities, the HumanEval Chen et al. (2021) and the MBPP Austin et al. (2021) are two widely used benchmarks. Additional problem-solving datasets include APPS Hendrycks et al. (2021), CodeContests Li et al. (2022), and xCodeEval Khan et al. (2023), among others.
Specialized Benchmarks:
There are also specialized benchmarks that measure specific capabilities of LLMs. For instance, the MT-Bench Zheng et al. (2024)) evaluates whether LLMs can properly engage in conversations, the RewardBench Lambert et al. (2024) assesses the performance of reward models. Other specialized benchmarks like the AlpacaEval777https://tatsu-lab.github.io/alpaca_eval/ evaluates the instruction following capabilities Zhou et al. (2023a) of LLMs, the Open Medical-LLM Leaderboard888https://huggingface.co/blog/leaderboard-medicalllm evaluates the biomedical capabilities of LLMs, HHEM999https://huggingface.co/spaces/vectara/leaderboard leaderboard for hallucination detection Mishra et al. (2024); Sadat et al. (2023), BigCodeBench101010https://huggingface.co/blog/leaderboard-bigcodebench and LiveCodeBench111111https://huggingface.co/blog/leaderboard-livecodebench for code generation capability evaluation, SWE-bench Jimenez et al. (2023) for software engineering capability evaluation. The recently proposed FOFO benchmark Xia et al. (2024) measures language models’ ability to adhere to the requested formats in prompts across different domains. Moreover, there are also some specialized benchmarks that are used for LLM safety121212https://huggingface.co/spaces/AI-Secure/llm-trustworthy-leaderboard Chao et al. (2024) and red teaming131313https://huggingface.co/spaces/HaizeLabs/red-teaming-resistance-benchmark Tedeschi et al. (2024) evaluation. The ability to understand both language and vision is another interesting capability of recently proposed LLMs Zhu et al. (2023a); Liu et al. (2024a, 2023b); Dai et al. (2024); Zhang et al. (2023); Ye et al. (2023b); Luo et al. (2024); Bai et al. (2023); Chen et al. (2023a). This has led to the development of many multi-modal benchmarks Fu et al. (2023a); Yu et al. (2023); Liu et al. (2023e); Li et al. (2023a, b); Liu et al. (2024a); Fu et al. (2024a); Lu et al. (2022); Guan et al. (2023); Li et al. (2023d); Qiu et al. (2024); Chen et al. (2024b). These benchmarks study the multimodal capabilities of LLMs across various domains, such as math and reasoning Yue et al. (2023); Lu et al. (2023), science diagrams Kembhavi et al. (2016), chart understanding and reasoning Masry et al. (2022); Islam et al. (2024b); Masry et al. (2024); Rahman et al. (2023), document understanding Mathew et al. (2021).
Other Diverse Benchmarks:
To enable a more comprehensive evaluation of LLMs across a wide range of scenarios, some studies also focused on introducing new benchmarks covering various aspects, such as HELM Liang et al. (2022), PromptBench Zhu et al. (2023b), OpenLLM141414https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, MixEval Ni et al. (2024), etc. These benchmarks cover diverse tasks and usually include existing benchmarking datasets (e.g., MMLU, HellaSwag, BoolQ Clark et al. (2019), etc.). Additionally, despite the availability of numerous benchmarks (both general and specialized), existing widely-used benchmarks still do not cover the full variety of tasks Preum et al. (2018); Parvez et al. (2018). Therefore, some researchers have independently evaluated LLMs using additional diverse datasets and tasks, including various NLP datasets and tasks Laskar et al. (2023a); Bang et al. (2023); Qin et al. (2023); Kocoń et al. (2023). They also employed domain-specific benchmarks in fields such as biomedicine Jahan et al. (2023, 2024), finance Li et al. (2023e); Guo et al. (2023a), language-specific Kabir et al. (2023); Lai et al. (2023); Khondaker et al. (2023); Ahuja et al. (2023); Liu et al. (2023c); Abdelali et al. (2024), social science Ziems et al. (2024), coding Liu et al. (2024b), and information retrieval Zhu et al. (2023c). In addition to that, ethics, bias, toxicity, robustness, and trustworthiness are also independently evaluated by researchers across various datasets Hendrycks et al. (2020a); Yang et al. (2022); Wang et al. (2023a); Zhuo et al. (2023); Rawte et al. (2023); Liu et al. (2023a); McIntosh et al. (2024); Sun et al. (2024).
A.2 Prominent LLMs
The impressive success of ChatGPT has led to the development of many LLMs in recent years. Since there are hundreds of LLMs being released in recent years Zhao et al. (2023a), we only discuss some of the prominent LLMs that achieved top rankings in various public leaderboards recently. LLMs can be categorized into two parts: Closed-Source LLMs: only available for use through the API or web interface, and (ii) Open-Source LLMs: where the pre-trained weights of the model are available that allow further training of such models. Below, we present some prominent LLMs in these two categories.
A.2.1 Closed Source LLMs
In the following, we categorize LLMs based on the organizations that develop these LLMs:
OpenAI models OpenAI (2023):
GPT-3.5: This model is an iteration of the GPT-3 architecture, emphasizing improvements in response quality through the application of the reinforcement learning from human feedback (RLHF) technique. GPT-3.5 is known for its robust performance in zero-shot tasks, where no specific training examples are provided during the task execution. This model has been instrumental due to its strong foundational capabilities in understanding and generating human-like text.
GPT-4: It extends GPT-3.5’s capabilities by incorporating multimodal functionalities, allowing the model to process not just text but also visual inputs. This advancement significantly broadens its applicational scope, making it adept at handling more complex tasks that require an understanding of both textual and visual information. It features enhanced safety protocols and a sophisticated training regime that includes a safety reward signal during its reinforcement learning phase.
GPT-4 Turbo: This version builds upon GPT-4’s foundation with substantial upgrades in computational efficiency and functionality. GPT-4 Turbo boasts an increased model capacity and an extended knowledge base that encompasses more recent data up to April 2023. It features a longer context window of up to 128,000 tokens and includes significant improvements in the model’s economy and output consistency.
GPT-4o: OpenAI’s most sophisticated model, GPT-4o ("o" for "omni") is a multimodal powerhouse capable of handling both text and image inputs to generate text outputs. It improves upon GPT-4 Turbo by offering double the text generation speed and reducing operational costs by 50%.
Google models:
PaLM-2: Released by Google in 2023, it is an advanced large language model that builds on the foundations set by its predecessor, the original PaLM. This iteration incorporates a sophisticated ’mixture of objectives’ technique, allowing it to surpass the capabilities of the earlier model significantly Anil et al. (2023).
Gemini: It is a multimodal model developed by google in December 2023, to understand and process a variety of information types, including text, images, audio, and video, seamlessly. Gemini’s architecture allows it to perform exceptionally across multiple platforms, from large-scale data centers to mobile devices, adapting efficiently to the needs of different applications. This model sets new benchmarks in AI with its ability to excel in tasks that require complex multimodal integrations Team et al. (2023).
Anthropic Models:
The Claude series models, developed by Anthropic, represent a series of advanced language models designed to enhance user interaction through natural language understanding and generation. Starting with the original Claude, which excelled in tasks like summarization and creative writing, each subsequent model—Claude Instant, Claude 2.0, and the Claude 3 family (Haiku, Sonnet, and Opus)—has introduced significant improvements in processing speed, reasoning capabilities, and multimodal functionality. These models have a variety of uses, from quick response generation in Claude Instant to sophisticated multimodal understanding in Claude 3 Opus, showcasing their versatility and advanced AI technology to meet different user and enterprise needs151515https://www.anthropic.com/news/claude-3-family. The latest model in the Claude-3 series is the Claude-3.5-Sonnet161616https://www.anthropic.com/news/claude-3-5-sonnet model.
A.2.2 Open Source LLMs
We similarly categorize the open-source LLMs based on the organizations that develop them:
Meta Models:
Llama: Launched in February 2023 by Meta AI, Llama was the first in the Llama series, showcasing strong performance on a range of natural language processing tasks. It competed well against larger models like GPT-3 with a smaller parameter size and was made available under a non-commercial license, primarily for academic research Touvron et al. (2023a).
Llama 2: Released in July 2023, Llama 2 improved on its predecessor by expanding model sizes up to 70 billion parameters. It maintained the original architecture but included better training data and enhanced functionality. Notably, Llama 2 was more accessible, available for both academic and some commercial uses Touvron et al. (2023b).
Llama 3: In April 2024, Meta AI introduced Llama 3171717https://llama.meta.com/llama3/, the most advanced version with up to 70 billion parameters. This version added longer context capabilities and improved multimodal functions, marking a significant advancement in AI technology application across various fields.
Mistral Models:
Mistral AI, founded in April 2023, is specialized in the development of open-source large language models. Rapidly gaining recognition in the AI industry, Mistral AI emphasizes the importance of open-source software, providing a viable alternative to proprietary models. The company has released several models, including Mistral 7B, Mixtral 8x7B, and Mixtral 8x22B, which are known for their high performance and innovation in the use of mixture of experts architectures Jiang et al. (2023). Codestral 22B, introduced on May 29, 2024, is a pioneering code generation model designed to enhance coding efficiency across more than 80 programming languages. With its specialized focus and lightweight architecture, Codestral significantly outperforms other leading models on the HumanEval FIM benchmark, making it a critical tool for developers seeking advanced AI-assisted coding capabilities.
Alibaba Models:
QWEN series models are transformer-based large language models developed by Alibaba Cloud Bai et al. (2023). These models, pre-trained on diverse data sources including web texts, books, code, and more, come in various sizes ranging from 0.5 billion to 110 billion parameters. Qwen models support long context lengths and demonstrate strong performance on multiple Chinese and English evaluation tasks, including common-sense reasoning, code, and mathematics. The latest versions, Qwen 1.5 and Qwen 2, offer significant improvements in chat model performance, multilingual support, and stable support for up to 32K context length. With a comprehensive vocabulary of over 150K tokens, Qwen models are designed to handle multiple languages effectively, making them a versatile tool for various AI applications.
Microsoft Models:
The Phi series Abdin et al. (2024) by Microsoft consists of small language models (SLMs) designed to provide high performance with lower computational requirements. The newly announced Phi-3 family includes models like Phi-3-mini, Phi-3-small, and Phi-3-medium, ranging from 3.8 billion to 14 billion parameters. These models excel in various benchmarks, offering capabilities similar to larger models but in a smaller, more cost-effective package. Phi-3 models are particularly suited for simpler tasks, local device operations, and environments with limited resources, making AI more accessible and efficient for diverse applications. They are available through Microsoft Azure AI Model Catalog, Hugging Face, and as NVIDIA NIM microservices. Several followup works extends Phi-models or their synthetic data into multilingual space such as Boughorbel et al. (2024).
Technology Innovation Institute Models:
Technology Innovation Institute release the Falcon series models Almazrouei et al. (2023), such as the Falcon 2 series that include models with parameter sizes of 1.3B, 7.5B, 40B, and 180B. These models are notable for their use of the REFINEDWEB dataset. Falcon models are designed for both research and commercial use, with Falcon 2 models featuring multilingual and multimodal capabilities, including vision-to-language. The Falcon 180B model, in particular, is accessible under a royalty-free license.
Cohere Models:
Cohere offers a variety of advanced large language models designed for multiple use cases, including text generation, embeddings, and reranking. The Command family models, such as Command R+ and Command R, excel in conversational tasks and complex workflows like code generation and retrieval-augmented generation (RAG) 181818https://cohere.com/command Lewis et al. (2020); Gao et al. (2023b); Chen et al. (2024a); Liu et al. (2023d); Xiong et al. (2024); Tang and Yang (2024); Alonso et al. (2024); Lyu et al. (2024); Parvez et al. (2021, 2023); Wang et al. (2023c). The Embed models enhance search, classification, and clustering capabilities with both English and multilingual support. The Rerank models improve search algorithms by re-organizing results based on specified parameters. Cohere models are accessible across platforms like Amazon SageMaker, Microsoft Azure, and Oracle GenAI Service, enabling seamless integration into diverse applications and retrieval augmented generation.
Google Gemma Models:
While early LLMs released by Google’s are mostly closed-source (e.g., PalM-2, Gemini, etc.), Google has also recently released some lightweight open-source LLMs, named as Gemma191919https://storage.googleapis.com/deepmind-media/gemma/gemma-report.pdf family LLMs, that also have multimodal capabilities202020https://huggingface.co/blog/paligemma.
A.3 Prompting Techniques
Prompts can be designed in various ways Schulhoff et al. (2024); Brown et al. (2020); Wei et al. (2022b); Chung et al. (2022); Islam et al. (2024a); Parvez (2024), as stated below:
In-Context Learning (Zero-shot): It means that the prompt used to interact with the model contains no examples or demonstrations. The model relies on its pre-existing knowledge, obtained from its initial training on diverse data, to generate a response or perform the task based solely on the instructions given. For example, “classify the sentence as biased or unbiased text”.
In-Context Learning (Few-shot): It means that the prompt used to interact with the model includes a small number of examples or demonstrations. The model uses these examples to quickly adapt and understand how to perform a specific task, leveraging the details within these examples. This technique allows the model to extend its pre-existing knowledge to new tasks by closely analyzing the limited examples given. For instance, classify the sentence as biased or unbiased based on a few similar examples provided.
Chain-of-Thought Prompting (CoT): This technique encourages models to generate intermediate reasoning steps before arriving at a final answer, mimicking a human-like problem-solving approach. This can be combined with few-shot prompting to achieve better results on more complex tasks. For example, if asked to determine whether the number "15" is odd or even, the model might outline its reasoning as follows: "An even number is divisible by 2 without a remainder. 15 divided by 2 is 7 with a remainder of 1. Therefore, 15 is an odd number." This step-by-step explanation helps clarify the model’s thought process and supports its conclusion.
Decomposition Techniques: These techniques break down complex problems into simpler sub-problems that can be solved sequentially by the GenAI model. Each component of the problem is addressed individually, and the solutions are integrated to form a comprehensive response. Decomposition is especially useful in tasks that require layered reasoning or have multiple steps. For example, in solving a math word problem, decomposition might involve separately calculating the distances each person travels and then combining these calculations to determine when they meet.
Role-based and Style-based Prompting: In these techniques prompts are designed to induce a specific style or persona in the model’s responses. By specifying a role (e.g., a scientist explaining a concept) or a style (e.g., formal or poetic), users can guide the tone and formality of the AI’s output. This technique is valuable in applications requiring genre-specific content generation or when the output needs to fit a particular communicative context.
Prompt chaining: It is a technique where a complex task is divided into simpler subtasks, each addressed by its own prompt. The response from one prompt is used as the input for the next, creating a sequential chain of prompts that gradually build towards the final answer. This method enhances the performance and reliability of large language models by breaking down tasks into manageable parts, making it easier to control and refine the model’s responses at each step. For example, in a document analysis task, the first prompt might extract key facts from a text, and the second prompt would use these facts to generate a summary.
Tree of Thoughts (ToT): It is a technique that structures problem-solving into a tree of possible solutions. It uses strategies like breadth-first or depth-first search to evaluate each potential solution path. For example, in solving a puzzle, ToT might explore different moves to find the quickest solution path.
Directional Stimulus Prompting (DSP) : It is a technique that enhances how large language models (LLMs) respond to tasks by using dynamically generated prompts. A secondary, tuneable model creates specific hints that guide the main, unchangeable LLM to produce more targeted and relevant outputs. This method uses reinforcement learning to refine these prompts based on how well they perform, making DSP a more adaptive and precise approach compared to standard prompting techniques. For instance, in summarizing complex documents, DSP might generate a prompt like "Summarize focusing on economic impacts," guiding the LLM to tailor its output specifically to the economic aspects mentioned in the text.
Multimodal Prompting: Extending beyond text, multimodal prompting involves using inputs like images, audio, or video along with textual descriptions. This technique leverages the model’s capability to process and integrate information from diverse data types, enhancing its applicability in scenarios where multiple forms of data are available. For example, interpret a scene from a video by analyzing both the spoken dialogue and the visual content to determine the mood of the conversation.
Meta-Prompting: It involves creating prompts that instruct the AI to generate or refine its prompts, essentially using AI to improve the efficiency and effectiveness of prompt engineering. This recursive use of prompting can lead to more dynamic and contextually adaptive AI behaviors. For example, ask the AI to optimize a prompt that instructs another AI to summarize news articles, thereby refining the instructions to enhance summary relevance and conciseness.
A.4 Decoding Parameters
There are various decoding parameters that are required to be set. For instance:
Temperature: It is used to control the randomness of the output. It is typically between 0 and 1. Lower values (e.g., 0.1) make the model more deterministic and focused on the most likely next token, while higher values (e.g., 0.9) introduce more randomness and diversity.
Beam Size: It refers to the number of beams in Beam Search Freitag and Al-Onaizan (2017), a decoding strategy that keeps track of multiple possible sequences (beams) at each step of generation to find the most likely sequence. A higher number of beams usually leads to more accurate results but at the cost of increased computation.
Top-K: The number of top probable tokens to consider. For example, if K=10, the model will choose the next token only from the top 10 most likely tokens.
Top-P: The cumulative probability threshold. For example, if P=0.9, the model will sample from the smallest set of tokens whose combined probability is at least 90%.
Maximum Output Tokens: It sets the maximum number of tokens to generate.
Scenario 1: For the response generated, designing a parsing script to extract the answer “Lionel Messi” is straightforward. However, the parsing script should also be robust to cover cases like abbreviations, uppercase-lowercase sensitivity, punctuations, synonyms, stemming, lemmatization, paraphrases, etc. |
Prompt: Which player has won the best player award in Fifa world cup 2022? |
Sample LLM Response (GPT 4o): Lionel Messi won the Best Player award (Golden Ball) in the FIFA World Cup 2022. He was instrumental in leading Argentina to victory in the tournament, culminating in their triumph in the final against France. |
Correct Answer: Lionel Messi |
Scenario 2: While Extraction of the answer “Lionel Messi” is required, due to the LLM knowledge-cut-off date of September 2021, it may answer about 2018. However, the target answer “Lionel Messi” is also in the output and so if the parsing script only parses the target answer then it may consider the response as correct whereas the response is wrong. |
Prompt: Which player has won the best player award in the last Fifa world cup? |
Sample LLM Response (Older ChatGPT 3.5 having knowledge cut-off date of September 2021): The Best Player award (Golden Ball) in the previous FIFA World Cup, which was held in 2018 in Russia, was won by Luka Modric from Croatia. Prior to the that, Lionel Messi had won it in 2014. |
Correct Answer: Lionel Messi |
A.5 Parsing Script Design
While there are various evaluation software Biderman et al. (2024); Dalvi et al. (2023) currently available, they are limited to certain scenarios (e.g., limited to certain datasets and benchmarks, prompts, etc.). Thus, for the evaluation of LLMs across diverse settings, researchers often require to write parsing scripts. We present some scenarios in Table 4 to demonstrate why parsing script is required for such cases and the importance of validating parsing scripts.
A.6 Evaluation Approach
A.6.1 Automatic Evaluation
To provide a high-level overview, automatic evaluation for LLMs can be divided into the following:
Language Modeling: Perplexity Jelinek et al. (1977) is widely used to study the performance of auto-regressive language models. It measures how confidently a model predicts the next word in a sequence, with the assumption that lower perplexity indicates better performance. Hence, perplexity has been historically used to assess the language model’s capability to generate a coherent language and is also useful to quickly compare different models or checkpoints.
Discriminative Tasks: For tasks involving class prediction, post-processing using a parsing script is usually required to extract answers from the LLM-generated responses to compare against gold labels. In this context, metrics such as Exact Match, Accuracy, Precision, Recall, F1, are usually utilized in discriminative tasks Laskar et al. (2023a); Qin et al. (2023); Bang et al. (2023).
Generative Tasks: For generative tasks such as summarization or machine translation, parsing scripts are usually not required Laskar et al. (2023a); Jahan et al. (2024) and so the full response generated by LLMs are compared against the gold reference. In this regard, ROUGE Lin (2004) and BLEU Papineni et al. (2002) which are based on n-gram word matching are widely used. Meanwhile, various contextualized similarity Parvez and Chang (2021); Laskar et al. (2020) metrics (e.g., BERTScore Zhang et al. (2019), AlignScore Zha et al. (2023); Wang et al. (2024a)) are also utilized that do not depend on word-based similarity measures.
A.6.2 Human Evaluation
Since LLMs generate human-like responses, it is often required to conduct qualitative evaluation of their responses. Earlier, qualitative evaluation of model-generated responses in terms of fluency, coherence, and informativeness were very popular Laskar et al. (2022b). However, with LLMs usually generating informative, fluent, and coherent response Laskar et al. (2023a); Bang et al. (2023); Qin et al. (2023); Kocoń et al. (2023), the evaluation of factual consistency of LLM-generated responses has become more important recently Fu et al. (2023b). Moreover, qualitative evaluation to compare between LLM-generated responses via leveraging humans based on the Elo rating system Zheng et al. (2024) has gained a lot of attention.
Elo Rating:
Elo rating works by comparing LLMs in pairwise “A vs B” comparisons, where each model is assigned an initial numerical rating Boubdir et al. (2023); Zhao et al. (2023b). The outcome of each comparison adjusts these ratings based on the Elo algorithm: if a model performs better than expected, its rating increases; if it performs worse, its rating decreases. The expectation of a model’s performance is calculated using its rating relative to its opponent’s, adjusted by a factor that represents the sensitivity of expected scores to differences in ratings. To ensure a robust evaluation of LLMs using the Elo benchmark, it’s important to follow key indicators like reliability and transitivity Boubdir et al. (2023). Reliability keeps Elo ratings consistent across various comparison sequences and prevents them from being overly sensitive to changes in hyperparameters, such as the K-factor. Transitivity is crucial, indicating that if model A is rated higher than model B, and model B is rated higher than model C, model A should logically rank above model C. Extensive testing with both synthetic and real-world data is essential to verify that Elo scores accurately and stably reflect model performance Boubdir et al. (2023). This involves making precise adjustments to the comparison order, selecting hyperparameters carefully, and utilizing numerous permutations to ensure outcome consistency. Due to the sensitive nature of the Elo rating system towards the order in which the updates were performed, Zheng et al. (2024) used the Bradley-Terry (BTL) model for their chatbot arena ranking. It is observed that model A can have a higher win rate than model B both empirically and statistically but a lower Elo rating. Since win rate serves as the stand-in measure for the probability of a model being better than another, this signifies the findings by Boubdir et al. (2023) that Elo rating is non-transitive with or without (BTL). On the other hand, BTL-based rating is tolerant to an imbalanced number of votes per model as shown by Zheng et al. (2024), they also propose a different probability of win rates that are derived from the ratings found from BTL which is transitive but doesn’t correlate with the empirical win rates.
Elo hacking: Crowdsourced Elo-based ranking has gained popularity through the LMSys leaderboard 212121https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard and has been accepted by various organizations, prompting them to release their LLMs early into this ecosystem for human evaluation. However, such setups can be easily exploited on a large scale using simple techniques. Figure 5 illustrates how someone can initially bypass the blind scoring mechanism through ownership hacking. Additionally, the evaluation of knowledge bases is not easily tracked, making votes on highly complex reasoning questions equivalent to those on simpler queries. Furthermore, upon the release of a popular model, systematic attacks or boosting can be initiated through ownership hacking. In addition to that, considering same score for tie and both-bad can significantly change leaderboard position. We recommend to use tie as point and both-bad as point.
A.6.3 LLMs as Evaluators
Since human evaluation is time-consuming Laskar et al. (2023c, d) and difficult to reproduce, the instruction-following capabilities of LLMs have also inspired researchers to use certain LLMs as a judge to evaluate the responses generated by other LLMs Perez et al. (2022); Huang et al. (2024a); Fu et al. (2023b); Hada et al. (2023); Chern et al. (2024); Lu et al. (2024); Kobayashi et al. (2024); Kocmi and Federmann (2023); Shankar et al. (2024); Luo et al. (2023); Gao et al. (2023a); Kim et al. (2024b). While prior work mostly utilized general-purpose closed-source LLMs-as-a-judge, the recently proposed Prometheus 2 Kim et al. (2024a) model is an open-source variant which is specifically trained for qualitative evaluation of model-generated responses and demonstrated higher correlation with humans.
However, research by Wang et al. (2023b) and Shen et al. (2023) has highlighted potential limitations in using LLM as evaluators, suggesting that while LLMs can excel in specific areas like translation quality and grammatical error correction ( Kobayashi et al. (2024); Kocmi and Federmann (2023)), their effectiveness as evaluators may vary significantly across different tasks. This highlights the ongoing debate and research into the capabilities and limitations of LLMs as evaluators in diverse linguistic domains.